OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 1 6x + 2y = −9 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (johnweldon1993):

Well lets solve for 'y' in that top equation....you'll see why 3x + y = 1 subtrac 3x from both sides to isolate 'y' y = -3x + 1 This is what we now know 'y' equals... Now lets look at that second equation 6x + 2y = -9 We know what 'y' equals...so lets replace 'y' with that 6x + 2(-3x + 1) = -9 *notice how the only variable left is 'x' now...can you solve this?

OpenStudy (anonymous):

ya distribute 2

OpenStudy (johnweldon1993):

Right...so we have 6x - 6x + 2 = -9 What next...?

OpenStudy (anonymous):

like terms 6 and -6

OpenStudy (johnweldon1993):

Which leaves us with 2 = -9 There is no 'x' left and these numbers do not match...so what is your answer...?

OpenStudy (anonymous):

no value for x

OpenStudy (johnweldon1993):

There you go!

OpenStudy (anonymous):

thanks dude

OpenStudy (johnweldon1993):

no problem